# Rapid Word Learning Through Meta In-Context Learning

### v2

@@ Appendix F Evaluation of Generated New Usage Examples
Tables 11 and 12 show samples of example sentences from the BabyLM-10M test portion and the Chimera dataset along with the new usage examples generated by both the Llama-3 8B baseline and the Minnow model finetuned from that baseline. Table 14 contains error analysis of new examples generated from the BabyLM-10M test portion, extending Table 2 in Section 5.3. For comparing the Minnow model finetuned from Llama-2 7B and the CoLLEGe model on Llama-2 7B, we provide examples generated from the Chimera dataset and the BabyLM-10M test portion in Table 13 and 15, respectively.
@@
$\bullet$ tales similar to " bluebeard " include " the fitcher’s bird " by the brothers [new-token] and the english tale " mr. fox ". $\bullet$ these types of tales include royalty and fantastical magic, many of which are published by brothers [new-token] and hans christian anderson. $\bullet$ previously on [new-token]$\bullet$ the house in the wood from the german of [new-token]. previously in [new-token] the story is actually entitled " the bridal of trier " as mentioned on brothers [new-token] site. Baseline wins. Minnow finetuned output simply changes the preposition of an example, which is likely incorrect. The baseline output copies “brothers [new-token]”, which is likely correct. grimm
@@ Appendix G Evaluation of Generated Definitions
Table 16 shows the full set of results of comparing the model-generated and ground-truth definitions from all models. Table 17 shows the average of 1-, 2-, and 3-shot results on the CoLLEGe-DefGen dataset. Tables 18 and 19 show additional definitions generated from the CoLLEGe-DefGen and Oxford test set by the baselines and the Minnow models (in addition to Table 4 in Section 5.4).
@@
Table 16: Quantitative evaluation of generated definitions by comparing them with ground-truth definitions. This table extends Table 3 in the main text by adding additional results of the Llama-2 7B baseline and FLAN-T5 models. The FLAN-T5 models generally perform better than all other models on the Oxford dataset, but note that the Oxford dataset is in-distribution for these models, and these models may be overfitting to this dataset (see Table 19 for examples and discussion).
@@
In their groundbreaking research, the team of geneticists successfully deactivated the [new-token], resulting in the unexpected bloom of dormant traits within the lab mice. a type of gene that codes for a protein, known as a repressor, which inhibits the expression of one or more genes by binding to the operator or associated silencers. a gene that is turned off in a cell a hypothetical new word that does not yet exist in the English language. repressor gene
@@
The contractor recommended [new-token] for insulation due to its excellent thermal resistance and fireproofing properties. a type of insulation material made from melted rock or industrial waste that is spun into a fibrous structure. It is used for thermal insulation, fire protection, and soundproofing in both residential and commercial buildings. a material used to insulate a building a unit of insulation used in construction mineral wool
@@ Appendix H Concepts of “Word”
In our off-the-shelf baselines (Section 3.4.1 and the additional specialized FLAN-T5 models in Section 5.4) and comparison of generations (Appendix E), we replace the word to learn by a pseudo-word, like “dax” or “wug”, regardless of the word’s syntactic category and other aspects of meaning. The pseudo-word is then tokenized, usually by a subword tokenizer for LLMs (thus may have multiple tokens). We choose the pseudo-word to be meaningless and commonly used in linguistic tests. However, a pre-trained LLM like Llama-3 may have priors of certain aspects of the pseudo-word’s meaning based on its form. One aspect of the meaning is syntax. For example, from the sentence “Susie goes skiing in the winter”, we replace “skiing” with “dax” and have the sentence “Susie goes dax in the winter.” The sentence has a problem: the part of speech of “skiing” is gerund, but “dax” does not look like a gerund (since it does not end in “-ing”). So the sentence could mislead an LLM like Llama-3, which can use morphological information from its subword tokenization. Another aspect of the meaning is semantics. For example, in Table 19, the baseline model sometimes sticks to its prior knowledge of the pseudo-word “wug,” as reflected in its generated definitions like “a made-up word” and “a type of bird” (“wug” referred to a bird-like creature in the Wug Test of Berko, 1958). We admit that this problem may weaken our baselines and comparison of generations. Future work should use more suitable pseudo-words, preserving the morphological inflections while removing the semantic information.
@@ Appendix I Changes in Other Capabilities
Table 20: Accuracies on BLiMP (Warstadt et al., 2020). We show the mean and the standard deviation (in the bracket) of 3 runs of Minnow. Minnow accuracies are very close to the pre-trained model.

### v3

@@ Appendix F Evaluation of Generated New Usage Examples
Tables 18 and 19 show samples of example sentences from the BabyLM-10M test portion and the Chimera dataset along with the new usage examples generated by both the Llama-3 8B baseline and the Minnow model finetuned from that baseline. Table 21 contains error analysis of new examples generated from the BabyLM-10M test portion, extending Table 2 in Section 5.3. For comparing the Minnow model finetuned from Llama-2 7B and the CoLLEGe model on Llama-2 7B, we provide examples generated from the Chimera dataset and the BabyLM-10M test portion in Table 20 and 22, respectively.
@@
[…]
@@ Appendix G Evaluation of Generated Definitions
Table 23 shows the full set of results of comparing the model-generated and ground-truth definitions from all models. Table 24 shows the average of 1-, 2-, and 3-shot results on the CoLLEGe-DefGen dataset. Tables 25 and 26 show additional definitions generated from the CoLLEGe-DefGen and Oxford test set by the baselines and the Minnow models (in addition to Table 5 in Section 5.4).
@@
Table 23: Quantitative evaluation of generated definitions by comparing them with ground-truth definitions. This table extends Table 4 in the main text by adding additional results of the Llama-2 7B baseline and FLAN-T5 models. No significant differences are found among Llama-2 7B models on BERTScore F1. The FLAN-T5 models generally perform better than all other models on the Oxford dataset, but note that the Oxford dataset is in-distribution for these models, and these models may be overfitting to this dataset (see Table 26 for examples and discussion).
@@
In their groundbreaking research, the team of geneticists successfully deactivated the [new-token], resulting in the unexpected bloom of dormant traits within the lab mice. a type of gene that codes for a protein, known as a repressor, which inhibits the expression of one or more genes by binding to the operator or associated silencers.
a gene that is turned off in a cell a hypothetical new word that does not yet exist in the English language. repressor gene
@@
The contractor recommended [new-token] for insulation due to its excellent thermal resistance and fireproofing properties. a type of insulation material made from melted rock or industrial waste that is spun into a fibrous structure. It is used for thermal insulation, fire protection, and soundproofing in both residential and commercial buildings.
a material used to insulate a building a unit of insulation used in construction mineral wool
@@ Appendix H Concepts of “Word”
In our off-the-shelf baselines (Section 3.4.1 and the additional specialized FLAN-T5 models in Section 5.4) and comparison of generations (Appendix E), we replace the word to learn by a pseudo-word, like “dax” or “wug”, regardless of the word’s syntactic category and other aspects of meaning. The pseudo-word is then tokenized, usually by a subword tokenizer for LLMs (thus may have multiple tokens). We choose the pseudo-word to be meaningless and commonly used in linguistic tests. However, a pre-trained LLM like Llama may have priors of certain aspects of the pseudo-word’s meaning based on its form. One aspect of the meaning is syntax. For example, from the sentence “Susie goes skiing in the winter”, we replace “skiing” with “dax” and have the sentence “Susie goes dax in the winter.” The sentence has a problem: the part of speech of “skiing” is gerund, but “dax” does not look like a gerund (since it does not end in “-ing”). So the sentence could mislead an LLM like Llama, which can use morphological information from its subword tokenization. Another aspect of the meaning is semantics. For example, in Table 26, the baseline model sometimes sticks to its prior knowledge of the pseudo-word “wug,” as reflected in its generated definitions like “a made-up word” and “a type of bird” (“wug” referred to a bird-like creature in the Wug Test of Berko, 1958). We admit that this problem may weaken our baselines and comparison of generations. Future work should use more suitable pseudo-words, preserving the morphological inflections while removing the semantic information.
@@ Appendix I Changes in Other Capabilities
Table 27: Accuracies on BLiMP (Warstadt et al., 2020). We show the mean and the standard deviation (in brackets) of 3 runs of Minnow. Minnow accuracies are very close to the pre-trained model.